# Finite Automata Extraction: Low-data World Model Learning as Programs from Gameplay Video

### v1

@@ Introduction
The programs were first learned using a training video, after which we used a test video for evaluation. To generate the frames from the model, we used 2 approaches. In Approach 1, we use the ground truth frame at each time step $t$ to predict the next frame with our model at time step $t+1$. In Approach 2, we only use the ground truth frame at time step $t=0$, after which the predicted frame at timestep $t$ is used to predict the frame at timestep $t+1$.
Domains
Figure 2: Screenshots of gameplay of both domains, Pac-Man and River Raid
@@
We tested our pipeline in two domains, Pac-Man and River Raid. Pac-Man is an arena-based game where the player navigates a level and collect pellets, which are scattered everywhere on the ground, while avoiding ghosts that can kill Pac-Man if they make contact. The game features 4 ghosts, Blinky, Inky, Pinky and Clyde. We used a simplified version of the game developed with Unity by removing all the ghosts except Inky. We also reduced the size of the map from 20x31 to 8x8. We recorded the gameplay videos for Pac-Man ourselves. We extracted frames at a rate of four frames per second as the average human reaction time is 0.25s (Thomas 1969). We specified a custom stationary background so that the loss could focus on the sprites. The frames were divided into batches of individual games. The width $w$ and height $h$ were set to 8, and the number of sprites in the dictionary $l$ was set to 15. We trained the model for 1000 epochs.
@@
River Raid is a scrolling shooter game where the player controls an airplane and has to shoot obstacles that come in its way. While recording, we cropped the screen to remove background details that were not involved in the gameplay to better learn the sprites. We used the Atari gym environment to record the gameplay videos for River Raid. We used the same hyper parameters from Pac-Man with the exception that $l$ was set to 16.
Baselines
For the world model, we used GameGAN (Kim et al. 2020) as our baseline. We believe that this model is a good representation of current world models, and models like GameNGen and Genie2 are not accessible to us. We first trained GameGan on the same PacMan data generated for our model to see if how it performs. .
We use Game Engine Learning (Guzdial, Li, and Riedl 2017b) as our baseline for our programs.
@@
IF (neighboring entities(PACMAN, POWERPELLET))
@@
Table 1: Programs Generated with Finite Automata Extractor for Pac-Man
@@
Table 2: Programs Generated with Finite Automata Extractor for River Raid
Metrics
We used FID and FVD, which are standard metrics for video generation (Liu et al. 2024) to assess the quality of the generated frames. In addition to using these established metrics, we created another metric, Prediction Error $\sum_{i=0}^{l}MAE(T_{i+1}-T_{i},M_{i+1}-T_{i})$ to assess the distance between the ground truth set of frames $T_{1},...T_{l}$} and generated set of frames $M_{1},...M_{l}$}. We compare this against our baseline, GameGan.
To assess the quality of the programs, we used the length of the programs and the number of conditions as the two metrics gives us a good assessment of the generalizability of the programs.
Results
In this section we first cover the learned programs for Pac-Man and River Raid domains and then cover results comparing the performance of FAE against the baselines within the two domain.
Learned Programs
We were able to learn six distinct sprites for Pac-Man, PACMAN, BLINKY, GHOST, EYES, POWERPELLET, EMPTY. PACMAN indicates the player, BLINKY indicates the red ghost sprite, GHOST indicates the blue ghost sprite that BLINKY turns into when PACMAN eats a POWERPELLET, EYES indicates the eye sprite a GHOST turns into when eaten, and EMPTY represents empty space. Notably these names are given only for reporting purposes and do not exist in the system. We could not learn a sprite for the normal pellets due to their size, which is a limitation of Marionette. We include a translated version of the learned programs in Table 1. The DSL encodes If and Then statements implicitly, as such we prepended the “IF” and “THEN” statements. We also translated from the sprite dictionary indices into human-readable names (e.g. “PACMAN”). However, the simplicity of this translation and the associated programs themselves indicates a strong potential for explainability.
We identified PACMAN and EMPTY as sprites that would not have learnable programs, PACMAN due to player control and BLANK due to a lack of dynamic information. The model successfully learned that BLINKY should always follow PACMAN, which can be seen in the first condition of BLINKY. It also predicts that BLINKY should change into GHOST when PACMAN collides or eats a POWERPELLET. The model learned that POWERPELLET should change into PACMAN once it has been eaten by PACMAN. This may seem odd, but has the desired effect, as POWERPELLET sprites should disappear when eaten by PACMAN, thus the POWERPELLET turning into the PACMAN that ate it has the same effect. The model could not replicate the ground truth behavior for GHOST and EYES, this is due to the randomness involved in the sprites, which currently is not captured with our current DSL. The GHOST does learn about being eaten by PACMAN, however, instead of turning into EYES like in the original game, it just disappears. These results show that our system has learned two of the four programs correctly. It should be noted that these two entities occupy considerably more screen time than the other two.
For River Raid, we also learned six distinct sprites, PLANE, FUEL, GREENBOAT, BLACKBOAT and PELLET. PLANE indicates the triangular sprite that the player controls, FUEL indicates the can sprite, GREENBOAT and BLACKBOAT represents the green and black boat sprites respectively. The translated programs for this domain can be seen in 2. Like Pac-Man, we identified PLANE and EMPTY as sprites that would not have learnable programs. The model successfully learned that PELLET should always be moving up. However, it also learned to change to EMPTY for arbitrary positions. The model did not learn code for BLACKBOAT, this was due to the fact that the appearance of the sprite was rare in the training video, much like EYES in Pac-Man. FUEL is an interesting case since the size of the sprite is much bigger. As a result, it alternated between occupying two and three grid cells for each frame, depending on its position. This led to the algorithm not being able to learn it’s movement behavior. However, it did learn the collision behavior of the original game as it changes to empty upon colliding with PELLET. The behavior for GREENBOAT could not be replicated, we believe this was for the same reason as that for BLACKBOAT. These results are less favorable than those of Pac-Man, but demonstrate the importance of having a good representation since we did not account for sprites with different sizes.
Results
Method FID(train) FID(test) Error(train) Error(test) Training Time (s) Conditions
Game GAN 10.17 10.46 35.31 40.64 7200.00 -
Game Engine Learning 0.25 0.39 10.46 6.91 2.01 73±22.55
Finite Automata Extraction 0.23 0.28 10.60 7.06 801.23 1.75±0.83
@@ Results
Table 3: Comparison of methods on FID, Predicted Error, training time, and average number of conditions for Pac-Man, Approach 1.
Method FID(train) FID(test) Error(train) Error(test) Training Time (s) Conditions
Game GAN 10.17 10.46 35.31 40.64 7200.00 -
Game Engine Learning 0.19 0.23 10.87 7.16 2.01 73±22.55
Finite Automata Extraction 0.28 0.29 10.37 7.56 801.23 1.75±0.83
@@
Table 4: Comparison of methods on FID, Predicted Error, training time, and average number of conditions for Pac-Man, Approach 2
Method FID(train) FID(test) Error(train) Error(test) Training Time (s) Conditions
Game GAN 58.62 58.81 143.62 144.2 7200.00 -
Game Engine Learning 0.37 0.32 82.42 82.85 4.68 125.42±19.58
Finite Automata Extraction 0.29 0.25 82.60 83.03 817.23 2.00±1.87
@@
Table 5: Comparison of methods on FID, Predicted Error, training time, and average number of conditions for River Raid, Approach 1
Method FID(train) FID(test) Error(train) Error(test) Training Time (s) Conditions
Game GAN 58.62 58.81 143.62 144.2 7200.00 -
Game Engine Learning 0.17 0.14 83.44 84.14 4.68 125.42±19.58
Finite Automata Extraction 0.14 0.13 83.58 84.04 817.23 2.00±1.87
@@
Table 6: Comparison of methods on FID, Predicted Error, training time, and average number of conditions for River raid, Approach 2
@@
We first go over the results of the metrics for Pac-Man. The results for the Aproach 1 and Approach 2 can be seen in 3 and 4, which shows the FID, Predicted Error, time taken to learn, and average number of conditions across the models. Since Game Gan does not produce any code, the average number of conditions metric does not apply to it. While running the Game Engine Learning System, the algorithm failed to converge due to not being able to make a perfect prediction and so we take the best performing program which follows the practices from prior work(Guzdial, Li, and Riedl 2017b). The average number of conditions were less for Finite Automata Extraction than Game Engine Learning. This shows that the code extracted by our system was more general as fewer conditions would lead to more cases of learned code running. The difference in the number of conditions is based on the way the different algorithms work, GEL starts with the maximum number of conditions then reduces whereas we add conditions as needed.
We trained the GameGan model for 2 hours, thus it had a constant training time. Due to its Neural nature, despite having the highest training time, it had the worst performance. While GEL has the lowest training time, it failed to converge as stated before.
@@
For Approach 1, across the baselines, FAE had the lowest FID for both the train and test video, though the results are comparable with GEL. This is due to the fact that GameGan unable to learn a non-random model from the amount of available training data. Further, while GEL may appear to have similar performance this is due to it only predicting the passed in prior frame exactly. This indicates GEL did not learn transferrable information to this test video from the training video. However, predicting the prior frame is a strong baseline itself based on prior work(Guzdial, Li, and Riedl 2017b). For the same reason, the Predictive Error for GEL and FAE is again comparable.
@@
For Approach 2, GEL and FAE’s results were again comparable, although GEL had a lower FID for both test and train. It is to be noted that GEL did not make any changes to the frame, so the sequence output of the model consists of the same image. On the other hand, since FAE was not able to learn the behavior of GHOST and EYES sprites correctly, the image sequence completely diverges as soon as one of the two sprites appear in the ground truth video. The Predicted Errors are inconclusive since they are lower for FAE for the train video and lower for GEL for the test video.
@@
For River Raid, the FID and Predicted Error results for GameGan were again the worst due to model’s inability to learn a non-random model. GEL again had the best training time but with the caveat that it was unable to converge. The number of conditions were the less of FAE than GameGan. Notably, while the FID scores for GEL and FAE were again comparable, FAE had lower FID for both the approaches for both the train and test video.
@@
In the original River Raid game, Fuels and boats are spawned off screen randomly. This was not possible for our current DSL since we would need a Game Manager program for that, which is currently outside the scope of the current version of Retro Coder. Additionally, as noted in the Pac-Man case, Retro Coder currently cannot capture.
Limitations

### v2

@@ 5. Evaluation
The programs were first learned using a single training video of each game, after which we used a test video for evaluation. To generate the frames from the model, we used 2 approaches: Next Frame and Total Simulation. For Next Frame we use the ground truth frame at each time step $t$ to predict the next frame with our model at time step $t+1$. For Total Simulation, we only use the ground truth frame at time step $t=0$, after which the predicted frame at timestep $t$ is used to predict the frame at timestep $t+1$. Next Frame allows us to evaluate local coherency while Total Simulation evaluates global coherency of the learned programs.
5.1. Domains
Figure 3. Screenshots of gameplay of both domains, Pac-Man and River Raid.
@@ 5.1. Domains
Pac-Man is an arena-based game where the player navigates a level and collect pellets while avoiding ghosts that can kill Pac-Man if they make contact. The game normally features 4 ghosts, Blinky, Inky, Pinky and Clyde. We used a simplified version of the game developed with Unity by removing all the ghosts except Blinky. We also reduced the size of the map from 20x31 to 8x8. We recorded the gameplay videos for Pac-Man ourselves. We extracted frames at a rate of four frames per second based on the average human reaction time is 0.25s (Thomas, 1969).
The width $w$ and height $h$ of the learned grid were set to 8, and the maximum number of sprites in the dictionary $l$ was set to 15. We trained the model for 1000 epochs.
@@
River Raid is a scrolling shooter game where the player controls an airplane and has to shoot obstacles. While recording, we cropped the screen to remove background details that were not involved in the gameplay to better focus on the sprites and their behaviours. We used the Atari gym environment to record the gameplay videos for River Raid. We used the same hyper parameters from Pac-Man with the exception that $l$ was set to 16.
5.2. Baselines
For a neural world model baseline, we used GameGAN (Kim et al., 2020). We believe that this model is a reasonable representation of current world models. Further the largest of modern world models like GameNGen and Genie 3 are not accessible to the public, and could not be retrained with our resources even if they were. We trained GameGAN on the same Pac-Man and River Raid data we used to train FAE. Notably this is far less data than GameGAN had in the original publication. Our goal was to see how well a neural world model could do with this limited amount of data.
We use Game Engine Learning (Guzdial et al., 2017a) as a DSL-based world model baseline. Game Engine Learning is an existing symbolic world model that uses a top down approach, we wanted to see how well our algorithm performs again GEL. We again used the same training data that we used for FAE.
@@ 5.2. Baselines
IF (neighboring entities(PACMAN, PELLET))
@@
Table 1. Programs Generated with Finite Automata Extractor for Pac-Man
Sprite Learned Program
BLINKY
IF (exists in map(PACMAN))
THEN follow entity(PACMAN)
IF (neighboring entities(PACMAN, PELLET))
THEN change to entity(GHOST)
GHOST
IF (is neighboring(PACMAN))
THEN change to entity(EYES)
PELLET
IF (is neighboring(PACMAN))
THEN change to entity(PACMAN)
EYES
IF (exists in map(PACMAN))
THEN follow target location(4,4)
Table 2. Ground Truth Programs
@@
Table 3. Programs Generated with Finite Automata Extractor for River Raid
Sprite Learned Program
PELLET
IF (exists in map(PLANE))
THEN follow direction(UP)
IF (is neighboring(BLACKBOAT))
THEN change to entity(EMPTY)
IF (is neighboring(GREENBOAT))
THEN change to entity(EMPTY)
BLACKBOAT
IF (exists in map(PLANE))
THEN follow direction(DOWN)
IF (is neighboring(PELLET))
THEN change to entity(EMPTY)
FUEL
IF (exists in map(PLANE))
THEN follow direction(DOWN)
IF (is neighboring(PELLET))
THEN change to entity(EMPTY)
GREENBOAT
IF (exists in map(PLANE))
THEN follow direction(DOWN)
IF (is neighboring(PELLET))
THEN change to entity(EMPTY)
Table 4. Ground Truth Programs for River Raid
Method FID(train) FID(test) Error(train) Error(test) Conditions
GameGAN 10.17 10.46 35.31 40.64 -
Game Engine Learning 0.25 0.39 10.46 6.91 73±22.55
Finite Automata Extraction 0.23 0.28 10.60 7.06 1.75±0.83
@@
Table 5. Comparison of methods on FID, Predicted Error and average number of conditions for Pac-Man, with the Next Frame approach.
Method FID(train) FID(test) Error(train) Error(test) Conditions
GameGAN 10.17 10.46 35.31 40.64 -
Game Engine Learning 0.19 0.23 10.87 7.16 73±22.55
Finite Automata Extraction 0.28 0.29 10.37 7.56 1.75±0.83
@@
Table 6. Comparison of methods on FID, Predicted Error and average number of conditions for Pac-Man, with the Total Simulation approach.
Method FID(train) FID(test) Error(train) Error(test) Conditions
GameGAN 58.62 58.81 143.62 144.2 -
Game Engine Learning 0.37 0.32 82.42 82.85 125.42±19.58
Finite Automata Extraction 0.29 0.25 82.60 83.03 2.00±1.87
@@
Table 7. Comparison of methods on FID, Predicted Error and average number of conditions for River Raid, with the Next Frame approach.
Method FID(train) FID(test) Error(train) Error(test) Conditions
GameGAN 58.62 58.81 143.62 144.2 -
Game Engine Learning 0.17 0.14 83.44 84.14 125.42±19.58
Finite Automata Extraction 0.14 0.13 83.58 84.04 2.00±1.87
@@
Table 8. Comparison of methods on FID, Predicted Error and average number of conditions for River raid, with the Total Simulation approach.
5.3. Metrics
For a reconstruction quality metric we used FID, which is a standard metric for video generation (Liu et al., 2024). This allows us to capture a notion of the perceptual quality of the learned world models. In addition to using this established metric, we created another metric, Prediction Error $\sum_{i=0}^{l}MAE(T_{i+1}-T_{i},M_{i+1}-T_{i})$ to assess the distance between the ground truth set of frames $T_{1},...T_{l}$} and generated set of frames $M_{1},...M_{l}$}. This allowed us to capture a notion of precision over the sprites instead of over pixels.
It is difficult to assess the quality of learned programs. We note that the above metrics do indirectly capture some notion of the functional quality of the learned programs, especially Prediction Error. Outside of this, we also report the number of conditions, as a way to roughly approximate the generalizability of the learned programs. We only report this for FAE and Game Engine Learning, as GameGAN learns no explicit programs.
6. Results
In this section we first cover the learned programs for the Pac-Man and River Raid domains. Part of the benefits of FAE is that we can easily extract the learned programs and present them in a readable format. This also allows us to discuss some particulars of the learned programs that may be non-obvious from the metrics-based analysis. We then cover our metric results comparing the performance of FAE against the baselines within both domains.
6.1. Learned Programs
Using our variant of Marionette, FAE was able to learn six distinct sprite-classes for Pac-Man: PACMAN, BLINKY, GHOST, EYES, PELLET, EMPTY. PACMAN indicates the player, BLINKY indicates the red ghost sprite, GHOST indicates the blue ghost sprite that BLINKY turns into when PACMAN eats a PELLET, EYES indicates the eye sprite a GHOST turns into when eaten, and EMPTY represents empty space. Notably these names are given only for reporting purposes and do not exist in the system. FAE represents these entities are indices in the learned sprite dictionary. We could not learn a sprite-class for the normal pellets due to their size, which is a limitation of Marionette. We include a human-readable version of the learned programs in Table 2 and Table 4 using these entity names. The DSL encodes If and Then statements implicitly, as such we prepended the “IF” and “THEN” statements. However, the simplicity of this translation and the associated programs themselves indicates a strong potential for explainability.
We also included the closest ground truth programs for the two domains in Table 2 and Table 4. Do note that there were certain behaviours that could not be captured in our DSL and therefore the programs. We cover these behaviours later in this section.
For Pac-Man, we identified PACMAN and EMPTY as Non-learnable sprite-classes, PACMAN due to player control and BLANK due to a lack of dynamic information.
Even the ground truth program does not fully capture the correct behaviour for GHOST, indicating an issue with our DSL. In the actual game, GHOST moves in random directions. However, the current version of our DSL cannot capture random behaviour. As such this was not possible for FAE to learn. FAE successfully learned that BLINKY should always follow PACMAN, which can be seen in the first condition of BLINKY. It also predicts that BLINKY should change into GHOST when PACMAN collides or eats a PELLET. The third learned condition in BLINKY never gets satisfied in the game as the GHOST and BLINKY sprites are never together in a frame. This condition was learned due to the grids not being accurate all the time. Occasionally, MarioNette would hallucinate and return a grid with some sprites that do not actually exist in the true frame. The model learned that PELLET should change into PACMAN once it has been eaten by PACMAN. This may seem odd, but has the desired effect, a PELLET sprites should disappear when eaten by PACMAN, thus the PELLET turning into the PACMAN that ate it has the same effect. The model could not replicate the ground truth behaviour for EYES. This is because there were not enough frames in the training video to capture it’s behaviour. The GHOST does learn about being eaten by PACMAN, however, instead of turning into EYES like in the original game, it just disappears. These results show that our system has learned two of the four programs correctly. It should be noted that these two sprite-classes occupy considerably more screen time than the other two. We think this is a reasonable performance from a single gameplay example and could meaningfully support many game development tasks. We provide an example in the Case Study section.
For River Raid, we also learned six sprite-classes, PLANE, FUEL, GREENBOAT, BLACKBOAT and PELLET. PLANE indicates the triangular sprite that the player controls, FUEL indicates the can sprite, GREENBOAT and BLACKBOAT represents the green and black boat sprites respectively. The translated programs for this domain can be seen in 4. Like Pac-Man, we identified PLANE and EMPTY as Non-learnable sprite-classes.
Figure 4. Screenshots of gameplay of Pellet Chaser
While the ground truth programs for the game are accurate, there is one mechanic that they cannot cover. In the original game, BLACKBOAT, FUEL and GREENBOAT get randomly spawned on the top row. However, our current DSL cannot capture this behaviour. The model successfully learned that PELLET should always be moving up. However, it also learned to change to EMPTY for arbitrary positions. The model did not learn any code for BLACKBOAT, which we believe was because the sprite-class appeared rarely in the training video, much like EYES in Pac-Man. FUEL is an interesting case since the size of the sprite is much bigger. As a result, it alternated between occupying two and three grid cells for each frame, depending on its position. This led to the algorithm not being able to learn it’s movement behaviour. However, it did learn the collision behaviour of the original game as it changes to empty upon colliding with PELLET. The behaviour for GREENBOAT could not be replicated, which we believe was due to the same reason as BLACKBOAT. These results are less favourable than those of Pac-Man, and demonstrate the importance of DSL choices, since we assumed all sprites would occupy a single grid cell in our DSL.
6.2. Comparison with Baselines
@@
In this subsection we discuss the results of our metrics-based comparison with our two baselines. We first go over the results of the metrics for Pac-Man. The results for the two variations can be seen in Tables 5 and 6. These show the FID, Predicted Error, and average number of conditions across the models. Since GameGan does not produce any code, we cannot measure the average number of conditions. While running the Game Engine Learning System, the algorithm failed to converge due to not being able to make a perfect prediction and so we take the best performing program which follows the practices from prior work (Guzdial et al., 2017a). The average number of conditions were fewer for Finite Automata Extraction than Game Engine Learning (GEL). This shows that the code extracted by our system was more general as fewer conditions would lead to more cases of learned code running. The difference in the number of conditions is based on the way the different algorithms work, GEL starts with the maximum number of conditions then reduces whereas we add conditions as needed.
We trained GameGAN to until convergence. Due to its neural network nature and the low amount of data, it had the worst performance.
@@
For the Next Frame approach, across the baselines, FAE had the lowest FID for both the train and test video, though the results are comparable with GEL. This is due to the fact that GameGAN was unable to learn a non-random model from the amount of available training data. Further, while GEL may appear to have similar performance this is due to it only predicting the passed in prior frame exactly. This indicates GEL did not learn transferrable information to this test video from the training video. However, predicting the prior frame is a strong baseline itself as established by prior work (Guzdial et al., 2017a). For the same reason, the Predictive Error for GEL and FAE is again comparable.
@@
For the Total Simulation approach, GEL and FAE’s results were again comparable, although GEL had a lower FID for both test and train. This is because GEL did not make any changes to the frame, so every frame output from the model consists of the same image. On the other hand, since FAE was not able to learn the behaviour of GHOST and EYES sprites correctly, the image sequence completely diverges as soon as one of the two sprites appear in the ground truth video. The Predicted Errors are inconclusive since they are lower for FAE for the train video and lower for GEL for the test video.
@@
For River Raid, the FID and Predicted Error results for GameGAN were again the worst due to model’s inability to learn a non-random model. The number of conditions were also again less for FAE than GameGAN. Notably, while the FID scores for GEL and FAE were again comparable, FAE had lower FID for both the approaches for both the train and test videos.
@@
In the original River Raid game, fuel and boats are spawned offscreen randomly. This was not possible for our current DSL since we would need to model the area outside of the game screen, which is currently not possible in the current version of Retro Coder.
7. Case Study
In this section, we show a use case for how game developers could use our pipeline. We take the Pac-man and River Raid world model programs that we have learned, and then experiment with variations of them by altering the code and the sprites. We first make changes to the learned programs so that they resemble the ground truth programs, we assume the developer would have made changes to correct the issues from the learned programs, additionally we think these changes are small enough to be reasonable.
Figure 5. Screenshots of gameplay of River Raid Power Up
7.1. Example 1: Pellet Chaser
We changed the code in Pac-Man for PELLET so that it always chases PACMAN. We did this by changing only the if statement in PELLET. This changed the dynamics of the game quite a bit, as the pellets now all seek Pac-Man as seen in Figure 1. This changes the gameplay from collecting into just avoiding the ghost for long enough until the pellets can catch up with Pac-Man. While we don’t think this is a particularly good game on its own, we do find it interesting, and it’s the kind of simple thought experiment that FAE could afford to game developers.
Sprite Learned Program
BLINKY
IF (exists in map(PACMAN))
THEN follow entity(PACMAN)
IF (neighboring entities(PACMAN, PELLET))
THEN change to entity(GHOST)
GHOST
IF (is neighboring(PACMAN))
THEN change to entity(EYES)
PELLET
IF (exists in map(PACMAN))
THEN follow entity(PACMAN)
EYES
IF (exists in map(PACMAN))
THEN follow target location(4,4)
Table 9. Program for Pellet Chaser
7.2. Example 2: River Raid Power Up
We changed the code in River Raid so that the boats get destroyed whenever PLANE touches FUEL. We did this by adding an if statement to both the boats, the condition being that if PLANE and FUEL are neighbouring each other, the boats change to EMPTY. While this maintains the core loop of the game, it changes the dynamics substantially as the player now has the option to collide with FUEL to destroy the obstacles essentially giving the player an area-of-effect (AOE) attack that did not exist in the original game. Again, this is likely not a better game than River Raid, but it allows for an interesting and fast exploration of a variation of the game.
Sprite Learned Program
PELLET
IF (exists in map(PLANE))
THEN follow direction(UP)
IF (is neighboring(BLACKBOAT))
THEN change to entity(EMPTY)
IF (is neighboring(GREENBOAT))
THEN change to entity(EMPTY)
BLACKBOAT
IF (exists in map(PLANE))
THEN follow direction(DOWN)
IF (is neighboring(PELLET))
THEN change to entity(EMPTY)
IF (neighboring entites(PLANE, FUEL))
THEN change to entity(EMPTY)
FUEL
IF (exists in map(PLANE))
THEN follow direction(DOWN)
IF (is neighboring(PELLET))
THEN change to entity(EMPTY)
GREENBOAT
IF (exists in map(PLANE))
THEN follow direction(DOWN)
IF (is neighboring(PELLET))
THEN change to entity(EMPTY)
IF (neighboring entites(PLANE, FUEL))
THEN change to entity(EMPTY)
Table 10. Program for Pellet Chaser
8. Limitations